OpenStudy (anonymous):

will fan/medal!! solve for y: 4x+y=18 2x+y=9

OpenStudy (solomonzelman):

4x+y=18 2x+y=9 (multiplied the second equation times -2) 4x+y=18 -4x-2y=-18 add the systems. tell me what you get, please.

OpenStudy (anonymous):

-1y=0

OpenStudy (anonymous):

I don't know where to go from there

OpenStudy (anonymous):

Would it be 0?

OpenStudy (solomonzelman):

you solved for y, it is zero:)

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

anytime!